OpenStudy (anonymous):

Suppose that a and b are integers, a ≡ 11 (mod 19), and b ≡ 3 (mod 19). Find the integer c with 0 ≤ c ≤ 18 such that c ≡ a3 + 4b3 (mod 19).

OpenStudy (perl):

is that supposed to be \[c =a^3 + 4b^3 (\mod19)\]

OpenStudy (anonymous):

ya i forgot

OpenStudy (perl):

if a = 11 mod 19, then a^3 = 11^3 mod 19 = 1331 mod 19 = 1 mod 19

OpenStudy (perl):

because 1331 = 19*70 + 1

OpenStudy (perl):

remainder of 1 upon dividing 11^3 by 19

OpenStudy (anonymous):

ok but what about the +4b^3(mod19) part

OpenStudy (perl):

if b= 3 mod 19, then b^3 = 3^3 mod 19 = 27 mod 19 = 8 mod 19 then 4b^3 = 4* 8 mod 19 = 32 mod 19 = 13 mod 19

OpenStudy (perl):

now you can add those two simplifed expressions

OpenStudy (anonymous):

so ok, this is what i should add? (1 mod 19)+(13 mod 19) right

OpenStudy (perl):

yes

OpenStudy (anonymous):

wouldn't the answer be 14 mod 19?

OpenStudy (perl):

or you can just say the integer is 14

OpenStudy (anonymous):

thanks alot @perl

OpenStudy (perl):

your welcome :)